OpenStudy (anonymous):

what is the slope and y-intercept form of y=3

OpenStudy (anonymous):

Hmmmm do you know the equation y = bx + c ? at least i think thats the correct equation... ^^

OpenStudy (anonymous):

i only know y=mx+b

OpenStudy (saifoo.khan):

its y=mx+c

OpenStudy (saifoo.khan):

or in other words: y = bx +c

OpenStudy (anonymous):

ya we only learned y=mx+b but anywayz.. idk how to do this

OpenStudy (anonymous):

ya (: i thought i was rite but if u act like u arent sure its less embarrassing when ur wrong tee hee!!!

OpenStudy (anonymous):

haha

OpenStudy (saifoo.khan):

hehe.